OpenStudy (anonymous):

The illumination from a light source is directly proportional to the strength of the source and inversely proportional to the square of the distance from the source . Two light sources of intensities I sub 1 and 1 sub 2 are d units apart. What point on the line segment joining to the two sources has the least illumination?

OpenStudy (phi):

they are asking for the spot between the 2 lamps that has the least light The total light at a spot between them is 

OpenStudy (phi):

take the derivative and set it = to zero and solve for x

OpenStudy (anonymous):

when i already got the x.. is it already done?

OpenStudy (anonymous):

if you get the value of x then its over i think so

OpenStudy (phi):

Let's use A= \(I_1\) and B= \(I_2\) (easier to type) take the derivative with respect to x of the equation that gives the total illumination that reaches a spot between the two lamps: \[ \frac{d}{dx}(\frac{A}{x^2} +\frac{B}{(d-x)^2} )=0\] use the power rule \[ dx^n= n x^{n-1} dx\] \[ \frac{d}{dx} (A x^{-2} + B(d-x)^{-2} )= 0\] you get \[ -2 A x^{-3} -2 B (d-x)^{-3}(-1) =0 \] or \[ \frac{A}{x^3}= \frac{B}{(d-x)^3} \] \[ (\frac{(d-x)}{x})^3= \frac{B}{A} \] take the cube root of both sides \[ \frac{(d-x)}{x}= \sqrt[3]{\frac{B}{A}}\] let c= cube root of B/A for convenience: d-x= cx x+cx= d x= d/(1+c)